# Energy-ordered resource stratification as an agnostic signature of life

### v1

@@ Abstract
$f_{i\alpha}(\vec{N},\vec{R})$ $=k_{i\alpha}R_{\alpha}(x,t).$ (4)
The parameter $Y_{\alpha}$ represents the energy content of resource $\alpha$ and $m_{i}$ represents the maintenance energy for species $i$. For simplicity, we assume that each biological species is a specialist and consumes only one resource.
If the system were well-mixed, the single “fittest” biological species using the most energy-dense resource would outcompete all others and take over the system. In our model, diffusion across space promotes the coexistence of species using resources with different energy content.
Results
Simulations of our model show that self-replication and ecology lead to spatial stratification of biological species and chemical resources (Fig. 2b–c). Starting from no stratification, with homogeneous species profiles and identical resource profiles, species and resource dynamics naturally converge towards a steady state where the biological species, and consequently resources, become stratified (Fig. 2b–c). The emergence of stratification can be understood as follows: the species using the most energetic resource grows the fastest near the source $x=0$. Because it grows fastest, this species most strongly antagonizes all others near $x=0$ (Fig. 2b, blue). As a result, less energetic resources are not consumed near $x=0$ and penetrate further (Fig. 2b, orange, green, red and purple). The species using the next most energetic resource (Fig. 2b, orange) then grows the fastest in the adjacent region, and similarly inhibits the growth of others (Fig. 2b, green). This process continues as inhibited species and unconsumed resources diffuse further away from the source at $x=0$.
The resulting pattern of resource profiles is similarly stratified, as resources with progressively decreasing energy content are depleted deeper and deeper away from the source $x=0$ (Fig. 2c). We refer to this spatial pattern as energy-ordered resource stratification.
To quantify the degree of energy-ordered stratification, we first define the penetration depth of each resource. Specifically, for each resource, we define its penetration depth as the width of the rectangle with the same height as the resource concentration at the source $x=0$, and with an area equal to that of the resource profile (quantified numerically; pictorial representation in Fig. 3a). At the penetration depth, the area of the rectangle which does not overlap with the resource profile has the same area as the remaining resource profile (Fig. 3a; shaded area). For each simulation, we measure the “stratification order parameter” as the magnitude of the correlation between the energy content and penetration depth of all simulated resources (Fig. 3b).
To test the robustness of energy-ordered stratification, we repeat 1,000 simulations of our model across a variety of randomly chosen conditions (see Methods). Throughout simulations, we systematically vary two key parameters: the self-replication parameter $\gamma$ and the density of ecological interactions $\rho$. In each case, we quantify the mean stratification order parameter across simulations. We find that self-replication and ecology are not only sufficient, but also necessary to generate energy-ordered stratification (Fig. 3c).
In the absence of self-replication ($\gamma=0$), the energy content of a resource has no bearing on its spatial profile, with penetration depth set by diffusion and consumption rate (not energy content). In the absence of ecology ($\rho=0$), species coexist with no spatial segregation; as a result, all resources are co-utilized to depletion, with energy-ordered stratification again failing to emerge. In the presence of both, energy-ordered stratification emerges robustly, with the stratification order parameter rapidly transitioning to $\approx 1$ as both $\gamma$ and $\rho$ increase beyond 0 (Fig. 3c; dark red region). Taken together, self-replication and ecology are both sufficient and necessary to generate self-organized stratification of resources in order of their usable energy content.
Figure 3: Both self-replication and antagonism are necessary for robust stratification. (a) Quantification of resource penetration depth: for each simulated resource profile (blue and red), we find the width of the rectangle with area equal to that of the resource profile and the same initial height. (b) Quantification of stratification order parameter: for all resource profiles obtained from one simulation, we compute the magnitude of the correlation between their penetration depths and energy content $Y_{\alpha}$ (shown is an example from a simulation with 5 profiles). (c) Heatmap of the stratification order parameter over multiple simulations, where we systematically varied the self-replication parameter $\gamma$ and the density of ecological interactions $\rho$. Stratification does not emerge in the absence of either self-replication or ecology. As $\gamma$ and $\rho$ both increase, stratification emerges robustly.
Discussion
So far, the search for chemical signatures of extraterrestrial life has been largely confined to looking for specific molecules believed to be associated with biotic processes (e.g., phosphine [18]). This approach is almost inevitably limited to forms of life closely resembling those on Earth. Until recently, such candidate biosignatures have been the only feasible approach, since our access to faraway worlds has been limited to spectroscopic analysis. However, the advent of missions allowing direct probing of Mars soil or the “sample return missions” such as Hayabusa [19] or OSIRIS-REx [20] has opened up new possibilities.
Here, we proposed a new kind of biosignature, namely observing stratified spatial profiles of chemical compounds in order of their energy content. We showed that the combined action of two key ingredients—self-replication and interspecies antagonism—naturally leads to such a self-organized stratification. This is because (1) self-replication produces an emergent correlation between the energetic yield of a resource and its rate of depletion; and (2) the interspecies antagonism then allows metabolic processes to become spatially segregated. We used a minimal model to argue that these ingredients are both sufficient and necessary. This mechanism does not depend on specific hypotheses about the chemical implementation of the organisms or their metabolism, but only on the plausible assumption that for a sufficiently evolved life form, usable chemical energy in a compound (i.e., its growth yield on this resource) correlates with its total chemical energy. Thus, we propose that energy-ordered resource stratification might serve as a robust agnostic biosignature.
Stratified structures can also emerge abiotically, e.g., sedimentation and calcification [21]. However, these abiotic mechanisms are driven by solubility or other chemical properties, and are not expected to be correlated with the total chemical energy in a compound. In contrast, biotically, such a correlation emerges naturally, as our model illustrates. While we intentionally kept the ingredients to a minimum, other biotic mechanisms, e.g. metabolic cross-feeding, could lead to similar outcomes (e.g., Winogradsky columns [15, 22]). Broadening the range of biotic mechanisms supporting energy-ordered stratification as a biosignature could potentially make our proposal even more robust.
Acknowledgements. We are grateful to A. Murugan, R. Braakman, P. Byrne, G. Fournier and S. Seager for helpful discussions. This work was supported in part by the NSF grant PHY-2310746. A.G. acknowledges support from the Ashok and Gita Vaish Junior Researcher Award, as well as the Government of India’s DBT Ramalingaswami Fellowship.
Methods
We simulated our model by numerically evolving equations (1)–(2) with the assumptions as in equations (3) and (4). All simulations were done on a 1D domain of length $x\in[0,L]$ where $L=100$ in arbitrary units, assuming no boundary flux Neumann boundary conditions for species, and assuming resources entered at $x=0$ at flux $K$ and had no flux at $x=L$. For all simulations, we set $D_{N}=10$, $D_{R}=20$, $m_{i}=m=0.1$ and $k_{i\alpha}=1$ for all species and resources. For each simulation with $M$ resources and $S=M$ species, we chose the energy content $Y_{\alpha}$ of each resource $\alpha$ randomly from a uniform distribution between 0. For competitive interaction strengths $\alpha_{ij}$ between species $i$ and $j$, we first picked a fraction $\rho$ of the interactions randomly, setting the rest to zero. For the picked interaction strengths, we chose them by randomly selecting a number from a normal distribution with mean 0.4 and standard deviation 0.1. All diagonal entries $\alpha_{ii}=0$ were left out of this procedure. For initial conditions, we always set homogeneous initial conditions for species, while choosing quadractically decaying profiles for resources, numerically obtained to satisfy the boundary conditions.

### v2

@@ Acknowledgements
$f_{i\alpha}(\vec{N},\vec{R})$ $=k_{i\alpha}R_{\alpha}(x,t)N_{i}(x,t),$ (10)
$\ell=0.5$ is a leakage parameter that controls what fraction of consumed resources are secreted as metabolic byproducts, and $D_{\alpha\beta}$ is a cross-feeding matrix that encodes which resources $\alpha$ can be produced from which others $\beta$ as byproducts. For simplicity, we assumed that $D$ mimicked a linear chain: consumption of resource 1 released resource 2; consumption of resource 2 released resource 3; and so on. Similar to the main text, all simulations were done on a 1D domain of length $x\in[0,L]$ where $L=100$ in arbitrary units, assuming no boundary flux Neumann boundary conditions for species. In contrast with the main text, we assumed that only resource 1 was externally supplied, entering the system at $x=0$ at flux $K$; all other resources were internally generated via cross-feeding. No resource could escape the system, resulting in no flux boundary conditions at $x=L$ for all resources; and at $x=0$ for all cross-fed resources. For all simulations, we set $D_{N}=10$, $D_{R}=20$, $m_{i}=m=0.1$ and $k_{i\alpha}=1$ for all species and resources. We separately verified that small variations in these parameters, which break their degeneracy, do not affect our conclusions (simulations not shown). For each simulation with $M=3$ resources and $S=M=3$ species, we chose the energy content $Y_{\alpha}$ of each resource $\alpha$ as $Y_{1}=1.8,Y_{2}=1.3,Y_{3}=1$. For competitive interaction strengths $A_{ij}$ between species $i$ and $j$, we picked them randomly from a normal distribution with mean 0.4 and standard deviation 0.1. All diagonal entries $A_{ii}=0$ were left out of this procedure. For initial conditions, we set homogeneous initial conditions for species, while choosing quartically decaying profiles for resources, numerically obtained to satisfy the boundary conditions. A representative example of the long-time species and resource profiles is plotted in Fig. S1. As in the original model, we observe that energy-ordered resource stratification emerges even when only one resource is provided externally, as metabolic byproducts always have lower energy content and can only be depleted once they have diffused sufficiently away from the species producing them.
Including concentration-dependent growth rates & ecological redundancy
In our original model, we assumed that the growth (self-replication) rate of all species increases linearly with local resource concentration. While simple, this assumption might not hold when the local resource concentration is large. Moreover, it is common for the same resource to be consumed by multiple strains, some of which grow better at higher resource concentrations, while others grow better at lower concentrations (allowing them to coexist). To test whether such resource-dependence affects our conclusions about energy-ordered resource stratification, we simulated a variant of our model with the same equations (1) and (2), but with per-capita growth rates $g_{i}$ having the commonly assumed saturating form
$g_{i}(\vec{N},\vec{R})=\sum_{\alpha=1}^{M}Y_{\alpha}\frac{k_{i\alpha}R_{\alpha}(x,t)}{K_{i\alpha}+R_{\alpha}(x,t)}-m_{i}.$ (11)
To simulate an example of the consequences of changing these model assumptions, we first took the simple case of 3 species on 2 resources (Fig. S2a): 2 of which (blue) grew by consuming the same high-energy content resource (blue triangle), while the remaining (orange) consumed the second lower-energy content resource (orange triangle). For the blue species using the same resource, we assumed a trade-off in their maximal consumption rate $k_{i\alpha}$ and half-saturating resource concentration $K_{i\alpha}$ (Fig. S2b). We simulated this model using the same parameters and initial conditions as the original model in the main text, and observed that the two resources stratified in decreasing order of their energy content (Fig. S2d). The species also stratified in space, but there was significant overlap between two species: one that grew better at lower concentrations of the high-energy resource (light blue) and the one that consumed the low-energy resource (orange; Fig. S2c). This suggested two statements: (1) that energy-ordered resource stratification was robust to the concentration-dependent growth assumed, and (2) that resource stratification was stronger than species stratification.
To systematically test these two statements, we repeated our simulations for a more general case of 8 randomly generated species growing on 5 resources. The 3 resources with highest energy content each had two species capable of consuming them, one better at high concentration and the other at low concentration. The last 2 resources could only be consumed by 1 species each, and grew well at low concentrations ($K_{i\alpha}=10^{-3})$. We quantified the resource stratification order parameter as in the main text, by computing the penetration depth of each resource (Fig. S2f). To quantify the depth of species, we computed the expectation value of their position, i.e., the mean position using their species abundance profiles (Fig. S2e). A representative example is shown in Fig. S2g–h. We observed that our two statements indeed generalized to more species and resources. Importantly, we concluded that resources were still stratified in decreasing order of their energy content. Further, resource stratification was slightly stronger (stratification order parameter 0.9) than species stratification (stratification 0.7).
Varying the degree of generalism
In our original model, each biological species could exclusively consume 1 resource. To test if we would still observe stratification if species were instead generalists, i.e., allowed to consume more than 1 resource, we simulated a variant of our model where we could vary the degree of generalism. In this variant, we simulated our original model (Eqns (1)–(4)), with the only change being that the resource consumption matrix $k_{i\alpha}$ was not a diagonal matrix any more. Instead, we set all off-diagonal entries — corresponding to species consuming all resources — to a value $\Delta$ indicating the degree of generalism. We kept the diagonal entries at 1 in accordance with the model in the main text. In the original model $\Delta=0$. For each value of $\Delta$ between 0 and 1 that we tested, we simulated 10 randomly generated ecosystems, in each case measuring the resource stratification order parameter. As we increased $\Delta$ and all species became generalists, we found that the stratification order parameter at first remained close to 1 (Fig. S3). This suggested that energy-ordered resource stratification was robust to including generalists. If $\Delta$ was large and comparable to the diagonal entries of the consumption matrix $k_{i\alpha}$, this would imply a situation where all species were statistically indistinguishable in terms of their preferred energy-providing resource. In this situation, we would not expect resource stratification, and this was indeed what we observed (Fig. S3). Indeed, we found that beyond a certain degree of generalism $\Delta\approx 0.4$, resources ceased to be stratified.
Figure S1: Energy-ordered resource stratification also emerges from cross-feeding — when only one resource is supplied and the others are internally generated as byproducts. Simulating a minimal model incorporating self-replication, ecology — via both antagonism and cross-feeding (see supplementary text for details) — shows that cross-feeding also supports the emergence of spatially stratified profiles of (b) species and (c) resources. Shown here is an example from a simulation for 3 species and 3 resources (triangles; the last, grey resource acts as a “waste” and is not shown). Only the blue resource is supplied externally. As the blue species consumes this resource, it secretes the red resource, which the red species consumes, secreting the green resource, and so on.
Figure S2: Energy-ordered resource stratification is robust to concentration-dependent growth & ecological redundancy. (a) Schematic showing 3 species (squiggles) growing on 2 resources (blue and orange triangles) for a variant of our model with concentration-dependent growth modeled via Monod, or Michaelis-Menten kinetics (see supplementary text for details). In this example, the two species that both consume the high-energy blue resource grow better at high and low concentrations of this resource, respectively; the orange species grows well at high concentrations of the low-energy orange resource. (b) Growth rate profiles for the two blue species in (a) as a function of the concentration of the high-energy resource (blue triangle). Both species show a trade-off in their ability to grow well at high versus low concentrations of the resource. (c–d) Simulations of the model variant in (a) and (b) still show that (d) the two resources stratify by energy content while (c) species may be less stratified (e.g., the light blue and orange species appear weakly stratified). This model variant breaks the competitive exclusion principle by allowing species to coexist on fewer resources via the assumed trade-off in concentration-dependent growth. (e) We quantifying the typical depth of species by measuring their average abundance-weighted position given by their simulated profile (example shown). (f) We quantify the resource penetration depth akin to the main text. (g–h) show scatter plots of the the species and resource depths with their corresponding energy content for another representative example with 8 species growing on 5 resources (see supplementary text for details); for species we use the energy content of the resource they consume. One of the 8 species simulated eventually went extinct. The correlation between these quantities provides a stratification order parameter, which shows that — even in this model variant — resources remain strongly stratified by energy content.
Figure S3: As species use multiple resources with decreasing specificity of preference, stratification decreases. Plot showing the resource stratification order parameter (measured as shown in Fig. 3a) as a function of the degree of generalism $\Delta$ in species for 10 randomly generated ecosystems with $S=5$ species and $M=5$ resources. In this variant of our model, we allow species to consume all resources—i.e., to become generalists—by introducing a fixed off-diagonal term $\Delta$ in the consumption matrix $k_{i\alpha}$ (see supplementary text for details). For small $\Delta$, energy-ordered stratification persists. As $\Delta$ increases, energy-ordered stratification signature eventually disappears due to the increasing similarity between all biological species. Note that for $\Delta>\frac{1}{M-1}=0.25$, species obtain more energy from other resources than the one they “specialize” on. Values plotted indicate the mean resource stratification order parameter over 10 randomly generated systems for each value of $\Delta$, while the shaded intervals indicate the standard error in the mean.